Solve the system of equations by substitution.
5X-2Y=-1
X+4Y=35

Answer by mananth(16949)

5X-2Y=-1.....................1
X+4Y=35
X=35-4Y
plug value of X in 1
5(35-4Y)-2Y=-1
175-20y-2y=-1
-22y=-1-175
-22y=-176
/-22
y=8
plug value of y in 1
5x-16=-1
5x=15
/5
x=3
...
m.ananth@hotmail.ca